The mean points obtained in an aptitude examination is 179 points with a variance of 225 . What is the probability that the mean of the sample would be less than 175 points if 77 exams are sampled? Round your answer to four decimal places.
Answer #1
for normal distribution z score =(X-μ)/σx
here mean=       μ= 179
std deviation   =σ= 15.000
sample size       =n= 77
std error=σ=σ/√n= 1.7094
probability =P(X<175)=(Z<(175-179)/1.709)=P(Z<-2.34)=0.0096